when linda added up all her scores and divided by 11 her average was 80. Her teacher tells her she can drop the lowest score of 50. what is her new average?
Answer by richard1234(7193)

Sum of Linda's 11 scores = 11*80 = 880

After the 50 is dropped, the sum of Linda's 10 scores is 830.

New average = 830/10 = 83